# Lower semicontinuity, Stoilow factorization and principal maps

### v2

@@ 1. Introduction
$K_{A}\equiv\left\{\begin{array}[]{cl}|A|^{n}/\det A&\mbox{ if }\det A>0,\\ 1&\mbox{ if }A=0,\\ +\infty&\mbox{ if }A\neq 0\mbox{ and }\det A\leq 0,\end{array}\right.$
where $|A|\equiv\max_{x\in\mathbb{S}^{n-1}}|Ax|$ and $|Ax|$ is the usual euclidean norm of $Ax\in\mathbb{R}^{n}$. When $\Omega$ is an open subset of $\mathbb{R}^{n}$ and $u\colon\Omega\to\mathbb{R}^{n}$ is locally of Sobolev class $\mathrm{W}^{1,1}$ in $\Omega$, denoted $u\in\mathrm{W}^{1,1}_{\operatorname{loc}}(\Omega,\mathbb{R}^{n})$, the distortion function $K_{u}(x)\equiv K_{\textup{D}u(x)}$ is an $\mathscr{L}^{n}$-almost everywhere defined extended real-valued function on $\Omega$. If the distortion $K_{u}<+\infty$ almost everywhere and the Jacobian $J_{u}\equiv\det\textup{D}u\in\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)$, then $u$ is called a mapping of finite distortion, and if furthermore $K_{u}\in\mathrm{L}^{\infty}(\Omega)$, then $u$ is a weakly quasiregular map. A striking result of Reshetnyak states that a quasiregular map, so a map $u\in\mathrm{W}^{1,n}_{\operatorname{loc}}(\Omega,\mathbb{R}^{n})$ with $K_{u}\in\mathrm{L}^{\infty}(\Omega)$, admits a representative that is continuous, open and discrete. In the two-dimensional case, $n=2$, which is our focus here, this result can be further refined via the so-called Stoilow factorization, which asserts that for an open subset $\Omega$ of $\mathbb{C}$ a map $f\colon\Omega\to\mathbb{C}$ is continuous, open and discrete if and only if $f=h\circ F$, where $F\colon\mathbb{C}\to\mathbb{C}$ is a homeomorphism and $h\colon F(\Omega)\to\mathbb{C}$ is holomorphic (see [43], and [35] for a modern exposition).
In [31] Tadeusz Iwaniec and Vladimir Šverák brought together ideas from Geometric Function Theory and Nonlinear Elasticity proving, again in two dimensions, that if $u\in\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)\equiv\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega,\mathbb{C})$ has distortion function $K_{u}\in\mathrm{L}^{1}(\Omega)$, then (the precise representative of) $u$ is continuous, open and discrete (see also [5] for various refinements). The aim of this note is to explain the role of Stoilow factorization, and in particular of the result of Iwaniec and Šverák, in an apparently unrelated context, namely in proving lower semicontinuity results in situations where the energy functionals do not satisfy the standard growth assumptions, such as in typical variational models of hyperelasticity or variational problems arising in Geometric Function Theory.
@@
First recall that when $\Omega$ is a bounded open subset of $\mathbb{R}^{n}$ and the exponent $p\geq 1$, then the $\mathrm{W}^{1,p}$-quasiconvexity of a functional $\mathbf{E}\colon\mathbb{R}^{n\times n}\to\mathbb{R}$, meaning that
${\bf E}(A)\leq\fint_{\Omega}{\bf E}(\textup{D}u(z))\,\mathrm{d}m(z),\quad\;{\rm for\,all}\;u\in A+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})\;{\rm and}\;A\in\mathbb{R}^{2\times 2},$
@@
It is easy to see that this equivalence persists if the functional $\mathbf{E}$ satisfies (1.2) but, instead of being non-negative, satisfies instead the condition
@@
However, when $\mathbf{E}$ has only the standard $p$-growth (1.2) for an exponent $p>1$, then the variational integral (1.1) can fail to be $\mathrm{W}^{1,p}$-swlsc, due to possible concentration effects at the boundary [38, 14, 36, 1] (this happens already for $\mathbf{E}(A)=\det A$, $A\in\mathbb{R}^{2\times 2}$, and for $n=p=2$, see [23]).
@@
On the other hand, as observed by Meyers [37], the result persists if the variational integral $\mathscr{E}$ is restricted to Dirichlet classes: if $g\in\mathrm{W}^{1,p}(\mathbb{R}^{n},\mathbb{R}^{n})$, then for a functional $\mathbf{E}$ satisfying (1.2), the variational integral $\mathscr{E}$ is swlsc on $g+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})$ if and only if $\mathbf{E}$ is $\mathrm{W}^{1,p}$-quasiconvex. We refer to [17] for further discussions and to [21] for some related results also linking quasiconvexity to coercivity of $\mathscr{E}$ in Dirichlet classes.
In the absence of the standard $p$-growth condition (1.2), even for the Dirichlet classes $g+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})$ the sufficiency of $\mathrm{W}^{1,p}$-quasiconvexity for swlsc of $\mathscr{E}$ becomes unclear (in fact, counterexamples exist if the functional $\mathbf{E}$ is allowed to be discontinuous and extended real-valued), but the necessity of $\mathrm{W}^{1,p}$-quasiconvexity for $\mathrm{W}^{1,p}$-swlsc remains true [14].
In this situation one introduces a strengthening of the $\mathrm{W}^{1,p}$-quasiconvexity condition, the so-called closed $\mathrm{W}^{1,p}$-quasiconvexity. This reinstates the sufficiency for $\mathrm{W}^{1,p}$-swlsc of $\mathscr{E}$ on $\mathrm{W}^{1,p}(\Omega,\mathbb{R}^{n})$, provided the functional $\mathbf{E}$ satisfies also (1.3). But then the necessity becomes unclear (in fact, when $\mathbf{E}$ is allowed to be lower semicontinuous and extended real-valued, the necessity of closed $\mathrm{W}^{1,p}$-quasiconvexity fails, see [34]).
@@
These issues become particularly relevant in dealing with variational problems where the effective domain of the energy functional is the set
(1.4) +$\mathbb{R}^{2\times 2}_{+}\equiv\bigl{(}\mathbb{R}^{2\times 2}\cap\{\det>0\}\bigr{)}\cup\{0\},$
or in other words, the set of matrices $A$ with finite distortion $K_{A}$. Such functionals arise naturally for instance in elasticity, where e.g. the classical second invariant, the sum of the singular values divided by $\det(A)$, is given by
(1.5) $I_{2}(A)=K_{A}+\frac{1}{K_{A}},\qquad A\in\mathbb{R}^{2\times 2}_{+}.$
@@
However, two central aspects of the theory in this context remain somewhat obscure. Firstly, it is not clear how ${\bf E}(A)$ should diverge when $\det(A)$ tends to 00, to have a feasible theory for the corresponding variational integral $\mathscr{E}$. Secondly, it seems desirable to rule out pathological deformations (for example homeomorphisms whose determinant is zero a.e or sequences of maps which converge to such maps [28, 24]). Thus it is not obvious what are the appropriate function spaces in which minimisers should belong to, while experiments do not help to clarify these issues as we are close to a regime where the elasticity theory breaks down. For these reasons, and since the general structure of quasiconvex functionals is far from clear, the functionals defined on +$\mathbb{R}^{n\times n}_{+}$ that have been succesfully used in the elasticity literature, have mostly been polyconvex. See for example, [13, 22, 26] and particularly [11, 12] for the setting and success of polyconvexity.
@@
Similar obstacles were faced in our recent study [4] of the local Burkholder functional $\mathbf{B}_{K}$, see (1.8) and Subsection 1.2 for the details. This functional lives on the $K$-quasiconformal well
(1.6) $Q_{2}(K)\equiv\{A\in\mathbb{R}^{2\times 2}:\,K_{A}\leq K\},$
@@
while outside the well it is +$\infty$+ ∞, $\mathbf{B}_{K}(A)=+\infty$ when $A\notin Q_{2}(K)$. Because of sign of $\mathbf{B}_{K}$ on $Q_{2}(K)$ it is easy to see that the closed $\mathrm{W}^{1,p}$-quasiconvexity of $\mathbf{B}_{K}$ is a necessary condition for $\mathrm{W}^{1,p}$-swlsc of the associated variational integral. On the other hand, the sufficiency is now non-obvious. This is the opposite situation of what happens for non-negative extended real-valued functionals.
Such viewpoints apply to all dimensions, but for a planar homeomorphism an application of the chain rule yields the particular relation
$\int_{\Omega}\!K_{f}(z)\,\mathrm{d}m(z)=\int_{f(\Omega)}\!|\textup{D}f^{-1}(w)|^{2}\,\mathrm{d}m(w),$
bridging mappings of integrable distortion with the bi-Sobolev maps. Recalling (1.5), this gives yet another argument in favour of studying the $\mathrm{W}^{1,2}$-mappings with $K\in\mathrm{L}^{1}(\Omega)$.
@@ 1.1. Principal maps
To deal with the issues on $\mathrm{W}^{1,p}$-swlsc described above, we next turn to the role of Stoilow factorization in the setting of lower semicontinuity. Here the key notion is that of a principal map:
@@
A principal map is an orientation preserving homeomorphism $f\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,1}_{\operatorname{loc}}(\mathbb{C})$, with $K_{f}\in\mathrm{L}^{1}(\mathbb{D})$, which is conformal outside the unit disc $\mathbb{D}$ and has the Laurent expansion
$f(z)=b_{0}z+\frac{b_{1}}{z}+\phi(z),\quad{\rm where}\quad|b_{1}|<|b_{0}|\quad{\rm and}\quad\phi(z)\equiv\sum_{j=2}^{\infty}\frac{b_{j}}{z^{j}}\;\mbox{ for }\,|z|>1.$
@@
Since $1/z=\bar{z}$ for $|z|=1$, we have $f(z)=A_{f}(z)+\phi(z)$ for $|z|=1$, where the linear map $A_{f}\in\mathbb{R}^{2\times 2}$ (that we identify with its matrix) has in the complex notation the form $A_{f}(z)\equiv b_{0}z+b_{1}\bar{z}$. In particular, $A_{f}$ has positive determinant. We also record that
$\fint_{\mathbb{D}}\!\textup{D}f(z)\,\mathrm{d}m(z)=A_{f}.$
@@
Notice that it is standard to normalize principal maps to have $b_{0}=1$. We shall refer to such maps as normalized principal maps. It is useful to recall that the class of all normalized $\mathrm{W}^{1,2}$-principal maps forms a normal family (see [5]); in fact, any sequence of such normalized principal maps admits a subsequence that converges uniformly on $\mathbb{C}$ to a normalized principal map. On the other hand, principal maps with the mere $\mathrm{W}^{1,1}_{\operatorname{loc}}$-regularity do not form a normal family; a sequence of such Sobolev homeomorphisms can have a BV-limit.
@@
We also recall that the inverse of a principal map lies in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ and it has finite distortion, c.f. [28].
@@
A functional ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is called $\mathrm{W}^{1,p}$-principal quasiconvex if it satisfies the Jensen inequality for all principal maps of class $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$. More precisely, given a principal map $f\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$, we require that
(1.7) ${\bf E}(A_{f})\leq\fint_{\mathbb{D}}\!{\bf E}(\textup{D}f(z))\,\mathrm{d}m(z).$

### v3

@@ 1. Introduction
[…]
where $|A|\equiv\max_{x\in\mathbb{S}^{n-1}}|Ax|$ is the operator norm and $|Ax|$ is the usual euclidean norm of $Ax\in\mathbb{R}^{n}$. When $\Omega$ is an open, connected, non-empty subset of $\mathbb{R}^{n}$ (henceforth called a domain) and $u\colon\Omega\to\mathbb{R}^{n}$ is locally of Sobolev class $\mathrm{W}^{1,1}$ in $\Omega$, denoted $u\in\mathrm{W}^{1,1}_{\operatorname{loc}}(\Omega,\mathbb{R}^{n})$, its distortion function $K_{u}(x)\equiv K_{\textup{D}u(x)}$ is an $\mathscr{L}^{n}$-almost everywhere defined extended real-valued function on $\Omega$. If the distortion $K_{u}<+\infty$ almost everywhere and the Jacobian $J_{u}\equiv J_{\textup{D}u}\equiv\det\textup{D}u\in\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)$, then $u$ is called a mapping of finite distortion, and if furthermore $K_{u}\in\mathrm{L}^{\infty}(\Omega)$, then $u$ is called a weakly quasiregular map. A striking result of Reshetnyak states that a quasiregular map, so a map $u\in\mathrm{W}^{1,n}_{\operatorname{loc}}(\Omega,\mathbb{R}^{n})$ with $K_{u}\in\mathrm{L}^{\infty}(\Omega)$, is either constant or admits a representative that is continuous, open and discrete. In the two-dimensional case, $n=2$, that is our focus here, this result is a consequence of the so-called Stoilow factorization. When $\Omega$ is a domain in $\mathbb{C}$ it asserts that a non-constant map $f\colon\Omega\to\mathbb{C}$ is continuous, open and discrete if and only if it admits the factorization $f=H\circ F$, where $F\colon\Omega\to\mathbb{C}$ is a homeomorphism and $H\colon F(\Omega)\to\mathbb{C}$ is holomorphic (see [80], and [61] for a modern exposition). In [56] Tadeusz Iwaniec and Vladimir Šverák brought together ideas from Geometric Function Theory and Nonlinear Elasticity proving, again in two dimensions, that if $u\in\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)\equiv\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega,\mathbb{C})$ is not a.e. constant and has distortion function $K_{u}\in\mathrm{L}^{1}(\Omega)$, then (the precise representative of) $u$ is continuous, open and discrete (see also [8] for various refinements). The aim of this note is to give a systematic presentation of the notion of principal quasiconvexity considered first in [5]. In particular we aim to explain how it, in conjunction with Stoilow factorization and its extension by Iwaniec and Šverák, allows to prove lower semicontinuity results in situations where the energy functionals are not polyconvex and do not satisfy the standard growth assumptions. Relevant examples include the typical variational models for isotropic planar hyperelasticity and the variational problems arising in geometric function theory.
@@
First recall [66, 19] that when $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ and the exponent $p\geq 1$, then the $\mathrm{W}^{1,p}$-quasiconvexity of a functional $\mathbf{E}\colon\mathbb{R}^{n\times n}\to\mathbb{R}$, meaning that
${\bf E}(A)\leq\fint_{\Omega}{\bf E}(\textup{D}u(x))\,\mathrm{d}x\quad\;{\rm for\,all}\;u\in A+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})\;{\rm and}\;A\in\mathbb{R}^{2\times 2},$
@@
It is easy to see that this equivalence persists if the functional $\mathbf{E}$ satisfies (1.2) but, instead of being non-negative, satisfies the weaker condition
@@
However, when a $\mathrm{W}^{1,p}$-quasiconvex functional $\mathbf{E}$ has only the standard $p$-growth (1.2) for an exponent $p>1$, then the variational integral (1.1) can fail to be $\mathrm{W}^{1,p}$-swlsc, due to possible concentration effects at the boundary [66, 19, 62, 1, 39]. For instance, when $n=p=2$ an example of this with $\mathbf{E}(A)=\det A$, $A\in\mathbb{R}^{2\times 2}$, can be found in [33]. On the other hand, as observed by Meyers [65], the $\mathrm{W}^{1,p}$-swslc result persists provided the variational integral $\mathscr{E}$ is restricted to Dirichlet classes: if $g\in\mathrm{W}^{1,p}(\mathbb{R}^{n},\mathbb{R}^{n})$, then for a functional $\mathbf{E}$ satisfying (1.2), the variational integral $\mathscr{E}$ is $\mathrm{W}^{1,p}$-swlsc on $g+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})$ if and only if $\mathbf{E}$ is $\mathrm{W}^{1,p}$-quasiconvex. We refer to [23] for further discussions and to [31] for some related results also linking quasiconvexity to coercivity of $\mathscr{E}$ in Dirichlet classes.
In the absence of the standard $p$-growth condition (1.2) the problem of $\mathrm{W}^{1,p}$-swlsc is much harder, little can be said in the general case and most results in the literature require some additional assumptions for their validity. In fact already the definition of the variational integral $\mathscr{E}[u]$ is unclear at this level of generality. When $\bf E$} is allowed to be valued in $\mathbb{R}\cup\{+\infty\}$, or perhaps even in $\mathbb{R}\cup\{\pm\infty\}$, then we have issues already for smooth and compactly supported maps $u$. The approach we shall follow here is in a sense the naive one, namely to use (1.1), but with the integral interpreted as an upper Lebesgue integral. We often refer to this as the pointwise definition of the variational integral. The other approach to the definition of the variational integral is to define it via a relaxation process, which in many situations is more natural and often means that the ensuing variational integral has better properties. This approach is often called the Lebesgue-Serrin-Marcellini definition and in the present context goes back to [63]. If we use the pointwise definition of the variational integral (1.1), then both the $\mathrm{W}^{1,p}$-quasiconvexity and the rank-one convexity conditions remain necessary for $\mathrm{W}^{1,p}$-swlsc, see [19]. However, neither is sufficient. The rank-one convexity does not even imply $\mathrm{W}^{1,\infty}$-quasiconvexity, regardless of growth conditions, by Šverák’s counter example [82, 83] except if the mapping is into the target $\mathbb{R}^{2}$, where the problem remains open, see also [43]. In this connection we refer to [37] and [71] for the particularity of $\mathbb{R}^{2\times 2}$ matrices for this problem and also some partial results. The sufficiency of $\mathrm{W}^{1,p}$-quasiconvexity for $\mathrm{W}^{1,p}$-swlsc of $\mathscr{E}$ on the Dirichlet class $g+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})$ is an open problem when the functional $\bf E$} is assumed continuous. Easy counterexamples exist if the functional $\mathbf{E}$ is allowed to be discontinuous and extended real-valued, see [19]. In this situation one introduces a strengthening of the $\mathrm{W}^{1,p}$-quasiconvexity condition, the so-called closed $\mathrm{W}^{1,p}$-quasiconvexity [76]. It is a sufficient condition for $\mathrm{W}^{1,p}$-swlsc of $\mathscr{E}$ on $\mathrm{W}^{1,p}(\Omega,\mathbb{R}^{n})$, provided the functional $\mathbf{E}$ satisfies also (1.3). However, the necessity becomes unclear, and in fact, when $\mathbf{E}$ is allowed to be lower semicontinuous and extended real-valued, the necessity of closed $\mathrm{W}^{1,p}$-quasiconvexity fails, see [59].
@@
These issues become particularly relevant in dealing with variational problems where the effective domain of the energy functional is contained in the set of matrices with finite distortion $\{K_{A}<+\infty\}$. We denote
(1.4) $\mathbb{R}^{n\times n}_{+}\equiv\mathbb{R}^{n\times n}\cap\{\det>0\},$
and record that $\{K_{A}<+\infty\}=\mathbb{R}^{n\times n}_{+}\cup\{0\}$. Such functionals arise naturally in, for instance, elasticity [15, 16, 17] and in geometric function theory [8], where a common example is the classical second invariant
(1.5) $\mathbf{I_{2}}(A)=K_{A}+\frac{1}{K_{A}},\qquad A\in\mathbb{R}^{2\times 2}_{+}\cup\{0\},$
@@
which is discussed at length in [8, Chapt. 21] and in [9, 64]. However, for applications to elasticity in particular, two central aspects of the theory in this context remain somewhat obscure. Firstly, it is not clear how ${\bf E}(A)$ should diverge when $\det(A)$ tends to 0, in order to have a feasible theory for the corresponding variational integral $\mathscr{E}$. Secondly, it seems desirable to rule out pathological deformations (for example homeomorphisms whose determinant is zero a.e or sequences of maps which converge to such maps [49, 36]). Thus it is not obvious what are the appropriate function spaces in which minimisers should be sought, and unfortunately it appears that experiments do not help to clarify these issues partly because we are close to a regime where the elasticity theory breaks down. Let us emphasize that variational problems on the classes of $\mathrm{W}^{1,p}_{\operatorname{loc}}$ homeomorphisms, lead for $p\geq 2$ naturally to the so-called topologically monotone maps introduced by Morrey, see [8] and [53, 55], whereas for $p<2$ a more complicated condition, the no crossing condition, introduced by De Philippis and Pratelli in [34] becomes relevant (see also [29] for the case $p=1$). For these reasons, and since the general structure of quasiconvex functionals is far from clear, the functionals defined on $\mathbb{R}^{n\times n}_{+}$ that have been successfully used in the elasticity literature, have mostly been polyconvex. See for example, [17, 18, 32, 42] and particularly [15, 16, 68, 48] for the setting and success of polyconvexity.
@@
Similar obstacles were faced in our recent study [5] of the local Burkholder functional $\mathbf{B}_{K}^{\operatorname{loc}}$, see (1.13) and (1.14) below for the definitions. This functional is real-valued on the $K$-quasiconformal well
(1.6) $Q_{2}(K)\equiv\bigl{\{}A\in\mathbb{R}^{2\times 2}:\,K_{A}\leq K\bigr{\}},$
@@
while outside the well it is $\infty$, $\mathbf{B}_{K}^{\operatorname{loc}}(A)=+\infty$ when $A\notin Q_{2}(K)$. Because $\mathbf{B}_{K}^{\operatorname{loc}}$ is non-positive and $p$-homogeneous on $Q_{2}(K)$ it is easy to see that the closed $\mathrm{W}^{1,p}$-quasiconvexity of $\mathbf{B}_{K}^{\operatorname{loc}}$ is a necessary condition for $\mathrm{W}^{1,p}$-swlsc of the associated variational integral. On the other hand, the sufficiency is now non-obvious (assuming the critical case where K=p/(p2)K=p/(p-2)). This is exactly the opposite situation of what happens for non-negative extended real-valued functionals.
For later reference we finish this subsection by recalling a rather pleasant property of $\mathrm{W}^{1,1}$-homeomorphisms with integrable distortion: If $f\colon\Omega\to\mathbb{C}$ is a homeomorphism (onto $f(\Omega)$) of class $\mathrm{W}^{1,1}_{\operatorname{loc}}(\Omega)$ with $K_{f}\in\mathrm{L}^{1}(\Omega)$, then its inverse $f^{-1}\colon f(\Omega)\to\Omega$ is of class $\mathrm{W}^{1,2}(f(\Omega))$ and
(1.7) $\int_{\Omega}\!K_{f}(z)\,\mathrm{d}m(z)=\int_{f(\Omega)}\!|\textup{D}f^{-1}(w)|^{2}\,\mathrm{d}m(w).$
See [8], [49].
@@ 1.1. Principal maps
In order to address some of the issues described above we next turn to the role of Stoilow factorization in the setting of lower semicontinuity. Here the key notion is that of a principal map, which has for a long time been implicitly used in geometric complex analysis, for instance in connection with planar quasiconformal mappings and the Beltrami equations; explicitly the term is used in [8]. We require a slightly more flexible definition:
@@
A principal map is an orientation preserving homeomorphism $f\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,1}_{\operatorname{loc}}(\mathbb{C})$, with integrable distortion $K_{f}$ on the open unit disk $\mathbb{D}$, which is conformal outside the closed unit disc $\overline{\mathbb{D}}$ with the Laurent expansion
$f(z)=b_{0}z+\sum_{j=1}^{\infty}\frac{b_{j}}{z^{j}}\mbox{ for }|z|>1,$
where $|b_{1}|<|b_{0}|$.
In connection with the principal map $f$ we often write
$A_{f}(z)\equiv b_{0}z+b_{1}\bar{z}\mbox{ for $z\in\mathbb{C}$, and }\phi_{f}(z)\equiv\sum_{j=2}^{\infty}\frac{b_{j}}{z^{j}}\mbox{ for }|z|\geq 1.$
@@
Here the linear map $A_{f}\in\mathbb{R}^{2\times 2}$ (that we identify with its matrix) has in particular positive determinant $|b_{0}|^{2}-|b_{1}|^{2}$, and the function $\phi_{f}$ extends continuously to all of $\overline{\mathbb{C}}$, defining a holomorphic map on $\mathbb{C}\setminus\overline{\mathbb{D}}$. Since $1/z=\bar{z}$ for $|z|=1$, we hereby have $f(z)=A_{f}(z)+\phi_{f}(z)$ for $|z|=1$. We also record that
(1.8) $\fint_{\mathbb{D}}\!\textup{D}f(z)\,\mathrm{d}m(z)=A_{f}.$
@@
Notice that it is standard to normalize principal maps to have $b_{0}=1$. Here we shall refer to such maps as normalized principal maps. It is useful to recall that the class of all normalized $\mathrm{W}^{1,2}$-principal maps forms a normal family (see [8, Theorems 20.1.6 and 20.2.3]); in fact, any $\mathrm{W}^{1,2}_{\operatorname{loc}}$-bounded sequence of such normalized principal maps admits a subsequence that converges uniformly on $\mathbb{C}$ to a normalized principal map. On the other hand, normalized principal maps with the mere $\mathrm{W}^{1,1}_{\operatorname{loc}}$-regularity do not form a normal family; a sequence of such Sobolev homeomorphisms can have a BV-limit (see [8, Chapt. 21]).
@@
We also recall that via (1.7) the inverse of a principal map lies in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$. If the principal map itself lies in $\mathrm{W}^{1,2}_{\operatorname{loc}}$, then the inverse has likewise an integrable distortion.
Our next goal is the definition of $\mathrm{W}^{1,p}$-principal quasiconvexity, which is instrumental for the lower semicontinuity results in this paper. It can be viewed as a refinement, adapted to the setting described above, of the notions of $\mathrm{W}^{1,p}$-quasiconvexity that was introduced by Ball and Murat in [19] and of closed $\mathrm{W}^{1,p}$-quasiconvexity that was introduced by Pedregal in [75]. The condition is motivated by the Stoilow factorization mentioned above.
For one of our key examples it is important that we allow the considered integrands to be signed, so before giving our definition of principal quasiconvexity let us for comparison and precision also review the corresponding definitions of $\mathrm{W}^{1,p}$-quasiconvexity and closed $\mathrm{W}^{1,p}$-quasiconvexity in this context.
A lower semicontinuous functional ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is said to be $\mathrm{W}^{1,p}$-quasiconvex at $A\in\mathbb{R}^{2\times 2}_{+}$ provided
(1.9) ${\bf E}(A)\leq\fint_{\mathbb{D}}^{\ast}\!{\bf E}(\textup{D}f(z))\,\mathrm{d}m(z)$
holds for all $f\in A+\mathrm{W}^{1,p}_{0}(\mathbb{D})$ [^1] The asterix on the integral signifies that the integral is intended as an upper Lebesgue integral, meaning that when both ${\bf E}(\textup{D}f)^{+}$ and ${\bf E}(\textup{D}f)^{-}$ integrate to $\infty$ over $\mathbb{D}$, then the integral is taken to be $\infty$. The interested reader can verify that this is equivalent to requiring, for each $k\in\mathbb{N}$, the validity of the inequality (1.9) when $\bf E$} is replaced by $\max\{{\bf E},-k\}$.. If this Jensen inequality holds for all $A\in\mathbb{R}^{2\times 2}_{+}$, then $\bf E$} is said to be $\mathrm{W}^{1,p}$-quasiconvex.
As observed in [19] this condition depends on the exponent $p$ and it refines Morrey’s original definition of quasiconvexity in [66] that in this terminology corresponds to $\mathrm{W}^{1,\infty}$-quasiconvexity. The definition is arguably natural, but it still has some deficiencies. For instance, as observed in [19], an extended real-valued integrand need not be rank-one convex. The next definition strengthens the notion of $\mathrm{W}^{1,p}$-quasiconvexity and does always imply rank-one convexity.
A lower semicontinuous integrand ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is closed $\mathrm{W}^{1,p}$-quasiconvex at $A\in\mathbb{R}^{2\times 2}_{+}$ provided
(1.10) ${\bf E}(A)\leq\int_{\mathbb{R}^{2\times 2}_{+}}^{\ast}\!{\bf E}\,\mathrm{d}\nu$
holds for all homogeneous $\mathrm{W}^{1,p}$ gradient Young measures $\nu$ with centre of mass $A$. If it holds for all $A\in\mathbb{R}^{2\times 2}_{+}$, then $\bf E$} is said to be closed $\mathrm{W}^{1,p}$-quasiconvex. We refer for the definitions and basic properties of Young measures to the monographs [67, 76, 77]. In particular, we will use the notation and terminology of gradient Young measures as described in our paper [5]. In fact, closed quasiconvexity provides an abstract solution to a closely related lower semicontinuity problem when the considered integrands are assumed bounded from below, see [59]. We refer to Section 2 and also [5] for further discussion of these quasiconvexity notions. The main drawback is that it is very difficult to verify whether a $\mathrm{W}^{1,p}$-quasiconvex integrand is closed quasiconvex unless we have control on the generating sequence of the gradient Young measures.
Now we are ready for our definition of principal quasiconvexity.
@@
A lower semicontinuous integrand ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is $\mathrm{W}^{1,p}$-principal quasiconvex at $A\in\mathbb{R}^{2\times 2}_{+}$ if the Jensen inequality
(1.11) ${\bf E}(A)\leq\fint_{\mathbb{D}}^{\ast}\!{\bf E}(\textup{D}f(z))\,\mathrm{d}m(z)$
@@
holds for all principal maps $f\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$ with $\fint_{\mathbb{D}}\!\textup{D}f(z)\,\mathrm{d}m(z)=A$. We say that $\bf E$} is $\mathrm{W}^{1,p}$-principal quasiconvex if the Jensen inequality (1.11) holds for all $A\in\mathbb{R}^{2\times 2}_{+}$ and for the corresponding $\mathrm{W}^{1,p}_{\operatorname{loc}}$-principal maps $f$.
We have stated an abstract definition depending on $p$ parallel to that of $\mathrm{W}^{1,p}$-quasiconvexity and closed $\mathrm{W}^{1,p}$-quasiconvexity, however our examples all are $\mathrm{W}^{1,1}$-principal quasiconvex (Notice that we are testing the Jensen inequality with gradients of homeomorphism). We also remark that allowing for extended real-valued integrands is a matter of convenience, but it also means that it is easy to construct pathological examples. For example, a $\mathrm{W}^{1,1}$-principal quasiconvex integrand need not be rank-one convex: take the integrand that is zero at the two matrices $\mathrm{diag}(1,1)$, $\mathrm{diag}(1,2)$ and $\infty$ elsewhere. It is unclear if similar examples exist when the functional is required to be continuous.
Often in our lower semicontinuity theorems it is convenient that the integrands $\bf E$} are defined on the full matrix space $\mathbb{R}^{2\times 2}$ and not merely on $\mathbb{R}^{2\times 2}_{+}$. We achieve this by simply declaring
(1.12) ${\bf E}(A)=+\infty\,\mbox{ on }\,\mathbb{R}^{2\times 2}\setminus\bigl{(}\mathbb{R}^{2\times 2}_{+}\cup\{0\}\bigr{)}$
and throughout the entire paper this will be our standing assumption for the considered integrands that might not be explicitly defined beyond $\mathbb{R}^{2\times 2}_{+}$. The reader will notice that we did not specify any value for $\bf E$} at $0\in\mathbb{R}^{2\times 2}$. This is on purpose because the 0 matrix is in this context special and it is convenient to have an ad hoc approach here. Sometimes we extend $\bf E$} to 0 by lower semicontinuity and sometimes we declare that it is $\infty$ or some other value there.